### v2

@@ 3.4 Convergence of physical solutions
$\sup_{s\in[0,\underline{\tau}_{v}]}|B_{v}^{n}(s)-B_{v}(s)|+|x_{v}^{n}-x_{v}|+\sum_{u\in N_{v}^{-}(G)}|c_{uv}^{n}-c_{uv}|<\frac{\varepsilon}{4}.$
Then, there exists $\Delta^{\prime}<\frac{1}{4}\Delta$ such that
$\sup_{s,t\in[0,\underline{\tau}_{v}],|s-t|\leq 2\Delta^{\prime}}|B_{v}(s)-B_{v}(t)|<\frac{\varepsilon}{4}$
for any $v\in B_{G}(V_{0},1).$ Furthermore, as $\lim_{n\to\infty}\underline{\tau}_{v}^{n}=\tau_{v}$, there exists $N_{3}\geq N_{2}$ such that, for all $n\geq N_{3}$ and all $v\in B_{G}(V_{0},1)$, $|\tau_{v}-\underline{\tau}_{v}^{n}|<\Delta^{\prime}$.
@@
$0\geq x_{v}^{n}+B_{v}^{n}(\underline{\tau}_{v}^{n})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\}}.$
@@
$\varepsilon$ $<\inf_{s\in[0,\underline{\tau}_{v}-\Delta]}\underline{X}_{v}(s)\leq\underline{X}_{v}(\tau_{v}+2\Delta^{\prime})=x_{v}+B_{v}(\tau_{v}+2\Delta^{\prime})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{\tau_{v}+2\Delta^{\prime}}\}}$
$\leq x_{v}^{n}+B_{v}^{n}(\underline{\tau}_{v}^{n})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\}}+|x_{v}-x_{v}^{n}|+|B_{v}(\tau_{v}+2\Delta^{\prime})-B_{v}(\underline{\tau}_{v}^{n})|+|B_{v}(\underline{\tau}_{v}^{n})-B_{v}^{n}(\underline{\tau}_{v}^{n})|$
$\quad+\sum_{u\in N_{G}^{-}(v)}|c_{uv}^{n}-c_{uv}|+\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\setminus\underline{D}_{\tau_{v}+2\Delta^{\prime}}\}}$
@@
As a result, we can iteratively extract a sequence $(v_{m})_{m\geq 0}$ with the property that $\underline{\tau}_{v_{m}}>\tau_{v_{m}}$ and that each $v_{m+1}$ is a child of $v_{m}$ in $\mathcal{T}(G_{n},v_{m})\subset\mathcal{T}(G_{n},v_{0})$. This is a contradiction as $\mathcal{T}(G_{n},v_{0})\subset V_{0}$, the latter of which being a finite subset of $V$. Therefore, we obtain $D_{t}=\underline{D}_{t}$, for any $t\geq 0$ and thus
$X_{v}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{t}\}}=\underline{X}_{v}(t),\quad v\in V.$
@@
Following Remark 3.2, the pathwise solution map $\varphi$ induces a map $\Phi$ acting on probability distributions such that $\mathcal{L}(G,X,D)=\Phi(\mathcal{L}(G,c,x,B))$, which is proved to be continuous. As a result, we have actually obtained the continuity of $\varphi$ on the set of configurations which drive $\delta$-robust systems for some $\delta>0$.
Proof of Theorem 1.9.
Suppose $(G_{n},c^{n},x^{n})$ converges in distribution to $(G,c,x)$ in $\mathcal{G}_{*}[\mathbb{R}]$, and $(B_{v}^{n})_{v\in G_{n}}$ are independent standard Brownian motions conditioned on $(G_{n},c^{n},x^{n})$. Then it follows from an argument similar to that in the proof of [28, Proposition 2.14] that $(G_{n},c^{n},x^{n},B^{n})$ converges in distribution to $(G,c,x,B)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$, where $(B_{v})_{v\in G}$ are independent standard Brownian motions conditioned on $(G,c,x)$. The desired convergence then follows from Theorem 3.19 together with Lemma 2.3. ∎
@@ 3.5 Convergence of empirical measures
Theorem 3.21.
If a sequence of finite marked graph $(G_{n},c^{n},x^{n},B^{n})$ converges in probability in the local weak sense to $(G,c,x,B)$ of $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$ as $n\to\infty$, where $(G,c,x,B)$ drives $\delta$-robust systems for some $\delta>0$ and $B_{v}$ satisfies the downward crossing property for any $v\in G$, then $(G_{n},X^{n},D^{n})$ converges in probability in the local weak sense to $(G,X,D)$, and the empirical measure $\mu_{n}$ converges in probability to $\mathcal{L}(X_{o},\tau_{o})$ in $\mathcal{P}(\mathcal{D}\times\mathbb{R})$, where $o$ is the root of $G$.
Proof.
@@
$\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{\mathcal{C}_{v}(G_{n},c^{n},x^{n},B^{n})}\to\mathcal{L}(G,c,x,B)\quad\text{in probability},$
where $\mathcal{C}_{v}(G_{n},c^{n},x^{n},B^{n})$ denotes the connected component of the marked graph $(G_{n},c^{n},x^{n},B^{n})$ rooted at $v$. Let $\Phi$ be the map that assigns the law of its physical solution to the law of input configuration, which is well-defined and continuous by Theorem 3.19 at distribution measures of configurations that drive $\delta$-robust systems and have a driving noise component satisfying the downward crossing property. Then,
$\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{C_{v}(G_{n},X^{n},D^{n})}$ $=\lim_{n\to\infty}\Phi\Bigg{(}\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{C_{v}(G_{n},c^{n},x^{n},B^{n})}\Bigg{)}$
$=\Phi\left(\mathcal{L}(G,c,x,B))\right)$
@@
The second claim follows from the above combined with Remark 2.6 and Corollary A.3. ∎
Proof of Theorem 1.10.
Suppose $(G_{n},c^{n},x^{n})$ converges in probability in the local weak sense to $(G,c,x)$ in $\mathcal{G}_{*}[\mathbb{R}]$, and $(B_{v}^{n})_{v\in G_{n}}$ are independent standard Brownian motions conditioned on $(G_{n},c^{n},x^{n})$. Then it follows from an argument similar to that in the proof of [28, Corollary 2.16] that $(G_{n},c^{n},x^{n},B^{n})$ converges in probability in the local weak sense to $(G,c,x,B)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$, where $(B_{v})_{v\in G}$ are independent standard Brownian motions conditioned on $(G,c,x)$. The desired convergence then follows from Theorem 3.21 together with Lemma 2.3. ∎
@@ 3.5.1 Systems on regular trees
As the general theory has been proved, we move on to study some simple examples to illustrate how the general result can be applied. We start with the example $G=(V,E)$ where $V=\mathbb{Z}$, $E=\{(i,i+1)\mathrel{\mathop{\ordinarycolon}}\,i\in\mathbb{Z}\}$ and $(c_{i,i+1})_{i\in\mathbb{Z}}$ are non-negative random variables. The graph $G$ models an infinitely long chain of banks with uni-directional exposure. We assume $(c_{i,i+1},x_{i},B_{i})_{i\in\mathbb{Z}}$ are i.i.d. across $i\in\mathbb{Z}$ and that $(G,c,x,B)$ drives $\delta$-robust systems for some $\delta>0$ (the latter condition is not restrictive, as the percolation threshold for $G$ is $1$). It follows from Theorem 1.10 that
@@
$X_{0}(t)=x_{0}+B_{0}(t)-c_{-1,0}\mathbf{1}_{\{\tau_{-1}\leq t\}},$
$\tau_{-1}$ is independent of $(c_{-1,0},x_{0},B_{0})$, and that $\mathcal{L}(\tau_{-1})=\mathcal{L}(\tau_{0})$, we see that the cumulative distribution function $F_{\tau_{0}}$ of $\tau_{0}$ is a fixed point of the map $\Psi\mathrel{\mathop{\ordinarycolon}}\mathcal{P}([0,\infty])\to\mathcal{P}([0,\infty])$ defined as (here we identify a probability measure on $0,\infty]$ with its CDF)
$\Psi[F]_{t}$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}\Big{[}\inf_{s\in[0,t]}\left(x_{0}+B_{0}(s)-c_{-1,0}\mathbf{1}_{\{\tau_{-1}\leq s\}}\right)\leq 0\Big{]}\quad\tau_{-1}\sim F,\quad\tau_{-1}\perp\!\!\!\perp(c_{-1,0},x_{0},B_{0})$
$=\int_{[0,\infty]}\mathbb{P}\Big{[}\inf_{s\in[0,t]}\left(x_{0}+B_{0}(s)-c_{-1,0}\mathbf{1}_{[r,\infty)}(s)\right)\leq 0\Big{]}\,\mathrm{d}F(r),\quad t\in[0,\infty].$ (3.3)
@@
Proof.
$\underline{F}\mathrel{\mathop{\ordinarycolon}}=\lim_{n\to\infty}F^{n}$
exists in the sense of weak convergence of CDFs and gives the minimal fixed point of $\Psi$. To interpret this iteration probabilistically, define the truncated systems $G_{n}=G|_{[-n,n]}$ with $V_{n}\mathrel{\mathop{\ordinarycolon}}=\{i\in\mathbb{Z}\mathrel{\mathop{\ordinarycolon}}-n\leq i\leq n\}$ and $E_{n}\mathrel{\mathop{\ordinarycolon}}=\{(i,i+1)\mathrel{\mathop{\ordinarycolon}}-n\leq i<n\}$. Let the roots of $G_{n}$ and $G$ be the vertex 0, and define $(c^{n},x^{n},B^{n})$ on $G_{n}$ by the corresponding restriction of $(c,x,B)$. Apparently, $(G_{n},c^{n},x^{n},B^{n})$ converges almost surely to $(G,c,x,B)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$ as $n\to\infty$, which, in particular, implies $\mathcal{L}(\tau_{0}^{n})\to\mathcal{L}(\tau_{0})$ as $n\to\infty$ by Theorem 3.19 and Corollary A.3. As the vertex n-n has no in-neighbors in $G_{n}$, we see that
$\mathbb{P}[\tau_{-n}^{n}\leq t]=\mathbb{P}\Big{[}\inf_{s\in[0,t]}(x_{-n}+B_{-n}(s))\leq 0\Big{]},$
that is, $F_{\tau_{-n}^{n}}=\Psi[F^{0}]$. Moreover, as
$X_{i+1}^{n}(t)=x_{i+1}+B_{i+1}(t)-c_{i,i+1}\mathbf{1}_{\{\tau_{i}^{n}\leq t\}}$
and that $\tau_{i}^{n}$ is independent of $(c_{i,i+1},x_{i+1},B_{i+1})$, the latter of which has the same distribution as $(c_{-1,0},x_{0},B_{0})$, it holds that $F_{\tau_{i+1}^{n}}=\Psi[F_{\tau_{i}^{n}}]$ for $i\geq-n$. As a result, we obtain the relation $F_{\tau_{0}^{n}}=\Psi^{n}[F_{\tau_{-n}^{n}}]=F^{n+1}$, and the claim is proved. ∎
The above argument can also be applied to the analysis of financial networks modeled by $\mathbb{T}_{k+1}$, the infinite $k+1$-regular tree, in which each vertex (bank) has $1$ out-neighbor (creditor) and $k$ in-neighbors (debtors) and $(c_{v},x_{v},B_{v})_{v\in\mathbb{T}_{k+1}}$ are i.i.d. across $v\in\mathbb{T}_{k+1}$. We give the following characterization of the distribution of the default time $\tau_{0}$ where 0 denotes the root of $\mathbb{T}_{k+1}$. The proof, which we omit for brevity, can be done by truncating the tree at depths $n\to\infty$ and conducting a bottom-up induction on the truncated trees.
@@
$\Psi_{k}[F]_{t}$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}\Big{[}\inf_{s\in[0,t]}\Big{(}x_{0}+B_{0}(s)-\sum_{i=1}^{k}c_{i,0}\mathbf{1}_{\{\tau_{i}\leq s\}}\Big{)}\leq 0\Big{]}$
$\quad\quad(\tau_{i})_{1\leq i\leq k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F,\quad\quad(\tau_{i})_{1\leq i\leq k}\perp\!\!\!\perp((c_{i,0})_{1\leq i\leq k},x_{0},B_{0})$
$=\int_{[0,\infty]^{k}}\mathbb{P}\Big{[}\inf_{s\in[0,t]}\Big{(}x_{0}+B_{0}(s)-\sum_{i=1}^{k}c_{i,0}\mathbf{1}_{[r_{i},\infty)}(s)\Big{)}\leq 0\Big{]}\,\mathrm{d}F_{1}(r_{1})\cdots\mathrm{d}F_{k}(r_{k}).$
@@ 3.5.2 A preliminary numerical experiment
We remark that Propositions 3.22 and 3.23 provide a straightforward method to numerically approximate the distribution of the default time of a representative bank via iterating the map $\Psi_{k}$ and evaluating the probabilities with Monte-Carlo simulations. We plot the cumulative distribution functions of the default times corresponding to $k\in\{1,2,5,10\}$ as well as several sample paths of the process $X(t)$. In the experiment, we take $x_{0}$ to be distributed as the absolute value of $\mathcal{N}(10,4^{2})$, $B_{0}$ the standard Brownian motion, and $c=\frac{10}{k}$ for a fair comparison (so that the total exposure is kept constant). The number of iteration is taken to be $1000$, the total time horizon is $50$, the number of time steps is $1000$, and the number of Monte-Carlo samples is $5000$. The Python code for generating the above plot is attached in the supplementary materials.
As demonstrated by the numerical results, the specific problem configuration lies in the regime in which increased connectedness lowers the fraction of defaults. It is worth pointing out that this result does not contradict the finding of Remark 3.12: the latter illustrates that default cascades tend to be contained from spreading into distant parts in sparsely connected networks, while the example here focuses on the overall network healthiness in a macroscopic manner; these represent two different perspectives on systemic risk. A comprehensive theoretical analysis of the complex interplay between the overall healthiness and the topological structure of financial networks remains challenging and represents an important direction for future research.

### v3

@@ 3.4 Convergence of physical solutions
$|x_{v}^{n}-x_{v}|+\sum_{u\in N_{v}^{-}(G)}|c_{uv}^{n}-c_{uv}|<\frac{\varepsilon}{4}.$
Since each $Z_{v}(\cdot)$ is right continuous with no upward jumps, there exists $\Delta^{\prime}<\frac{1}{4}\Delta$ such that
$\sup_{s,t\in[0,\tau_{v}+\Delta],\,t\leq s\leq t+2\Delta^{\prime}}(Z_{v}(s)-Z_{v}(t))_{+}\leq\frac{\varepsilon}{4}$
for any $v\in B_{G}(V_{0},1)$ such that $\tau_{v}<\infty$. Furthermore, as $\lim_{n\to\infty}\underline{\tau}_{v}^{n}=\tau_{v}$, by Lemma B.5 and the assumption that $Z_{v}(\cdot)$ has no upward jumps, there exists $N_{3}\geq N_{2}$ such that, for all $n\geq N_{3}$ and all $v\in B_{G}(V_{0},1)$, $|\tau_{v}-\underline{\tau}_{v}^{n}|<\Delta^{\prime}$ and
$(Z_{v}(\underline{\tau}_{v}^{n})-Z_{v}^{n}(\underline{\tau}_{v}^{n}))_{+}\leq\frac{\varepsilon}{4}.$
@@
$0\geq x_{v}^{n}+Z_{v}^{n}(\underline{\tau}_{v}^{n})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\}}.$
@@
$\varepsilon$ $<\inf_{s\in[0,\underline{\tau}_{v}-\Delta]}\underline{X}_{v}(s)\leq\underline{X}_{v}(\tau_{v}+2\Delta^{\prime})=x_{v}+Z_{v}(\tau_{v}+2\Delta^{\prime})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{\tau_{v}+2\Delta^{\prime}}\}}$
$\leq x_{v}^{n}+Z_{v}^{n}(\underline{\tau}_{v}^{n})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\}}+|x_{v}-x_{v}^{n}|+(Z_{v}(\tau_{v}+2\Delta^{\prime})-Z_{v}(\underline{\tau}_{v}^{n}))_{+}$
$+(Z_{v}(\underline{\tau}_{v}^{n})-Z_{v}^{n}(\underline{\tau}_{v}^{n}))_{+}+\sum_{u\in N_{G}^{-}(v)}|c_{uv}^{n}-c_{uv}|+\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{\underline{\tau}_{v}^{n}}^{n(k_{v}^{n}-1)}\setminus\underline{D}_{\tau_{v}+2\Delta^{\prime}}\}}$
@@
As a result, we can iteratively extract a sequence $(v_{m})_{m\geq 0}$ with the property that $\underline{\tau}_{v_{m}}>\tau_{v_{m}}$ and that each $v_{m+1}$ is a child of $v_{m}$ in $\mathcal{T}(G_{n},v_{m})\subset\mathcal{T}(G_{n},v_{0})$. This is a contradiction as $\mathcal{T}(G_{n},v_{0})\subset V_{0}$, the latter of which being a finite subset of $V$. Therefore, we obtain $D_{t}=\underline{D}_{t}$, for any $t\geq 0$ and thus for $v\in V$,
$X_{v}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\underline{D}_{t}\}}=\underline{X}_{v}(t).$
@@
Following Remark 3.2, the pathwise solution map $\varphi$ induces a map $\Phi$ acting on probability distributions such that $\mathcal{L}(G,X,D)=\Phi(\mathcal{L}(G,c,x,Z))$, which is proved to be continuous. As a result, we have actually obtained the continuity of $\varphi$ on the set of configurations which drive $\delta$-robust systems for some $\delta>0$.
@@ 3.5 Convergence of empirical measures
Proof of Theorem 1.12.
@@
$\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{\mathcal{C}_{v}(G_{n},c^{n},x^{n},Z^{n})}\to\mathcal{L}(G,c,x,Z)\quad\text{in probability},$
where $\mathcal{C}_{v}(G_{n},c^{n},x^{n},Z^{n})$ denotes the connected component of the marked graph $(G_{n},c^{n},x^{n},Z^{n})$ rooted at $v$. Let $\Phi$ be the map that assigns the law of its physical solution to the law of input configuration, which is well-defined and continuous by Theorem 1.11 at distribution measures of configurations that drive $\delta$-robust systems and have a driving noise component satisfying the downward crossing property. Then,
$\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{C_{v}(G_{n},X^{n},D^{n})}$ $=\lim_{n\to\infty}\Phi\Bigg(\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{C_{v}(G_{n},c^{n},x^{n},Z^{n})}\Bigg)$
$=\Phi\left(\mathcal{L}(G,c,x,Z))\right)$
@@
The second claim follows from the above combined with Remark 2.6 and Corollary B.3. ∎
@@ 3.5.1 Systems on regular trees
As the general theory has been proved, we move on to study some simple examples to illustrate how the general result can be applied. We start with the example $G=(V,E)$ where $V=\mathbb{Z}$, $E=\{(i,i+1)\mathrel{\mathop{\ordinarycolon}}\,i\in\mathbb{Z}\}$ and $(c_{i,i+1})_{i\in\mathbb{Z}}$ are non-negative random variables. The graph $G$ models an infinitely long chain of banks with uni-directional exposure. We assume $(c_{i,i+1},x_{i},Z_{i})_{i\in\mathbb{Z}}$ are i.i.d. across $i\in\mathbb{Z}$ and that $(G,c,x,Z)$ drives $\delta$-robust systems for some $\delta>0$ (the latter condition is not restrictive, as the percolation threshold for $G$ is $1$). It follows from Theorem 1.12 that
@@
$X_{0}(t)=x_{0}+Z_{0}(t)-c_{-1,0}\mathbf{1}_{\{\tau_{-1}\leq t\}},$
$\tau_{-1}$ is independent of $(c_{-1,0},x_{0},Z_{0})$, and that $\mathcal{L}(\tau_{-1})=\mathcal{L}(\tau_{0})$, we see that the cumulative distribution function $F_{\tau_{0}}$ of $\tau_{0}$ is a fixed point of the map $\Psi\mathrel{\mathop{\ordinarycolon}}\mathcal{P}([0,\infty])\to\mathcal{P}([0,\infty])$ defined as (here we identify a probability measure on $0,\infty]$ with its CDF)
$\Psi[F]_{t}$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}\Big[\inf_{s\in[0,t]}\left(x_{0}+Z_{0}(s)-c_{-1,0}\mathbf{1}_{\{\tau_{-1}\leq s\}}\right)\leq 0\Big]\quad\tau_{-1}\sim F,\quad\tau_{-1}\perp\!\!\!\perp(c_{-1,0},x_{0},Z_{0})$ (3.4)
$=\int_{[0,\infty]}\mathbb{P}\Big[\inf_{s\in[0,t]}\left(x_{0}+Z_{0}(s)-c_{-1,0}\mathbf{1}_{[r,\infty)}(s)\right)\leq 0\Big]\,\mathrm{d}F(r),\quad t\in[0,\infty].$
@@
The proof will be given in Appendix C.
The same argument can also be applied to the analysis of financial networks modeled by $\mathbb{T}_{k+1}$, the infinite $k+1$-regular tree, in which each vertex (bank) has $1$ out-neighbor (creditor) and $k$ in-neighbors (debtors) and $(c_{v},x_{v},Z_{v})_{v\in\mathbb{T}_{k+1}}$ are i.i.d. across $v\in\mathbb{T}_{k+1}$. We give the following characterization of the distribution of the default time $\tau_{0}$ where 0 denotes the root of $\mathbb{T}_{k+1}$. The proof, which we omit for brevity, can be done by truncating the tree at depths $n\to\infty$ and conducting a bottom-up induction on the truncated trees.
@@
$\Psi_{k}[F]_{t}$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}\Big[\inf_{s\in[0,t]}\Big(x_{0}+Z_{0}(s)-\sum_{i=1}^{k}c_{i,0}\mathbf{1}_{\{\tau_{i}\leq s\}}\Big)\leq 0\Big]$
$\quad\quad(\tau_{i})_{1\leq i\leq k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F,\quad\quad(\tau_{i})_{1\leq i\leq k}\perp\!\!\!\perp((c_{i,0})_{1\leq i\leq k},x_{0},Z_{0})$
$=\int_{[0,\infty]^{k}}\mathbb{P}\Big[\inf_{s\in[0,t]}\Big(x_{0}+Z_{0}(s)-\sum_{i=1}^{k}c_{i,0}\mathbf{1}_{[r_{i},\infty)}(s)\Big)\leq 0\Big]\,\mathrm{d}F_{1}(r_{1})\cdots\mathrm{d}F_{k}(r_{k}).$
@@ 3.5.2 A preliminary numerical experiment
We remark that Propositions 3.16 and 3.17 provide a straightforward method to numerically approximate the distribution of the default time of a representative bank via iterating the map $\Psi_{k}$ and evaluating the probabilities with Monte-Carlo simulations. We plot the cumulative distribution functions of the default times corresponding to $k\in\{1,2,5,10\}$ as well as several sample paths of the process $X(t)$. In the experiment, we take $x_{0}$ to be distributed as the absolute value of $\mathcal{N}(10,4^{2})$, $Z_{0}$ the standard Brownian motion, and $c=\frac{10}{k}$ for a fair comparison (so that the total exposure is kept constant). The number of iteration is taken to be $1000$, the total time horizon is $50$, the number of time steps is $1000$, and the number of Monte-Carlo samples is $5000$. The Python code for generating the above plot is attached in the supplementary materials.
As demonstrated by the numerical results, the specific problem configuration lies in the regime in which increased connectedness lowers the fraction of defaults. A comprehensive theoretical analysis of the complex interplay between the overall health and the topological structure of financial networks remains challenging and represents an important direction for future research.